OpenStudy (prince1342):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) - 5? The graph shifts 5 units up. The graph shifts 5 units down. The graph shifts 5 units left. The graph shifts 5 units right.

OpenStudy (prince1342):

@Heated help

OpenStudy (prince1342):

@Bacon. @agent0smith @austinL @adrynicoleb @AngelCriner @BrenFC @Bewface @charlotte123 help me

OpenStudy (agent0smith):

If you graphed y = x^2 and then y = x^2 - 5, what would the second look like compared to the first? f(x) + k k shifts the graph up or down f(x-h) h shifts the graph left or right

OpenStudy (prince1342):

wut

OpenStudy (prince1342):

B?

OpenStudy (agent0smith):

f(x) + k k shifts the graph up or down f(x-h) h shifts the graph left or right Use those to help identify what f(x) - 5 does

OpenStudy (prince1342):

oh

OpenStudy (secret-ninja):

So B, right?